THE PERMANENT ELECTRIC DIPOLE MOMENT OF 3D-MONOHYDRIDES: CrD^{a}.

\underline{JINHAI CHEN}, TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

A description of the bonding in the ground and low-lying electronic states of the first-row transition metal monohydrides becomes complicated because of the simultaneous involvement of the 3d^{n} 4s^{2}, 3d^{n+1} 4s^{1} and, to some extent, 3d^{n} 4s^{1} 4p^{1} electron configurations of the metal. The permanent electric dipole moment can provide insight into this complicated bonding. Here we report on the determination of μ_{e} for chromium CrD. A number of low-N lines of the X^{6}Σ − A^{6}Σ (0,0) band of chromium monodeuteride, CrD, have been recorded at near the natural line width limit by combining a supersonic molecular beam with high resolution laser excitation spectroscopy. The analysis gives the permanent electric dipole moment of the X^{6}Σ(v=0) state as 3.51(3)D and 1.152(5)D for the A^{6}Σ(v=0) state can be measured with higher precision because of some interesting near degeneracies in its level structure. A comparison with theoretical predictions for CrH^{b}’^{c}’^{d}’^{e} and experimental values other 3d transition metal monohydrides will be given.

aSupported by DoE-Basic Energy Sciences

bD.P.Chong, S.R.Langhoff, C.W.Bauschlicher, Jr., S.P.Walch and H. Partridge, J. Chem. Phys. (85), 2850(1986).

cD.Dai, K. Balasubramanian, J. Mol. Spectrosc. 161, 455(1993).

dS.Koseki, T. Matsushita, and M.S. Gordon, J. Phys. Chem, 110, 2560(2006).

eG.Ghigo, B.O.Roos, P.C.Stancil, and P.F. Weck, J. Chem. Phys. 121, 8194(2004).